# Primeval very low-mass stars and brown dwarfs – VIII. The first age benchmark L subdwarf, a wide companion to a halo white dwarf

### v2

@@ 2 Wide binary selection
4757030391786232576 | GALEX J052933.2635653 | 71.2 | WD | (1) | |||
4757030327366948608 | 2MASS J052940266357091 | 47.92 | 70.8 | M8a | (2) | ||
3004346292322852096 | 2MASS J061004031031036 | 100.8 | WD | (3) | |||
3004346257963119872 | 2MASS J061000781031393 | 59.95 | 93.3 | M8.5b | (4) | ||
5863122429179888000 | VVV J125644.42620208.1 | 75.6 | DC WD | (4,5) | |||
5863122429178232704 | VVV J125641.09620203.8 | 18.20 | 70.4 | sdL3 | (6) | ||
1321738565727229056 | GALEX J155516.9+315307 | 59.8 | DA WD | (7) | |||
1321738561431758592 | SDSS J155517.36+315316.8 | 10.86 | 59.1 | sdM9.5a | (4) | ||
2056440344314369280 | IPHAS J202533.21+351509.6 | 91.1 | WD | (3) | |||
2056440275586819456 | 2MASS J20253351+3515035 | 7.40 | 92.2 | M8.5b | (4) |
@@
References: (1) Gaia DR3 (Gaia Collaboration et al., 2023) (2) Gentile Fusillo et al. (2021b); (3) This paper; (4) DECaPS DR2 (Schlafly et al., 2018; Saydjari et al., 2023); (5) Vista VVV (Minniti et al., 2010); (6) CatWISE2020 (Marocco et al., 2021); (7) SAND atmosphere models (Alvarado et al., 2024);
@@ 4.4 The age of cool white dwarf primary
The WD companion, VVV 125662A is located at the bottom right of the WD cooling sequence in the H-R diagram (Fig. 1), consistent with being an old, low-mass WD (Raddi et al., 2022). By fitting the Gaia EDR3 photometry of this WD, Gentile Fusillo et al. (2021a) estimated a K and , inferring a mass of about 0.52 M⊙. Gaia DR3 blue photometer (BP) and red photometer (RP) spectra (De Angeli et al., 2023; Montegriffo et al., 2023) are available for this source, but are rather noisy as expected for a faint object.
@@
The best-fits for our hydrogen-dominated and hydrogen-deficient models (Fig. 9) are numerically consistent, with reduced values close to unity. The hydrogen-dominated atmosphere model may be more realistic given the apparent decline in the near-infrared photometry that is not correctly captured in the models due to current uncertainties in the treatment of the H2–He collision-induced-absorption in cool WD atmospheres (Blouin et al., 2018). Our inferred atmosphere parameters of K and dex for the hydrogen-dominated model, or K and dex for the hydrogen-deficient model, yield similar masses of M⊙, and these results are compatible with those obtained by Gentile Fusillo et al. (2021a).
@@ 4.5 Kinematics
We integrated the Galactic orbit of VVV 125662AB assuming a static three-component potential that is representative of the Milky Way’s baryonic and dark matter components (the MWPotential2014
of the galpy
module for python
; Bovy (2015) for details). We accounted for the uncertainties of the velocity components be examining both the median trajectory and 1 variants. A visual representation of the Galactic orbit and the time evolution of the cylindrical radial and vertical coordinates are shown in Fig.,11. As reported in Primeval V, this orbit is highly eccentric (), prograde ( kpc km/s), and moving radially away from the Galactic centre.
@@ 4.6 Stability of the wide binary
For VVV 125662AB, = 1.2 pc or au. Hence, this system’s projected separation of 1375 au at a distance of 75.6 pc corresponds to 0.0052 , well within the limit at which the Galactic tidal field is unable to disrupt the system. It is therefore logical that VVV 125662AB has survived as a bound wide binary over its 10 Gyr lifetime.
@@ 5 Conclusions
VVV 125662A is a cool, likely hydrogen-dominated WD whose total age (cooling and progenitor) of 10.5 Gyr is also consistent with halo membership. Combined, these measurements yield a consistent picture of the VVV 125662AB system as a metal-poor, wide halo binary and a unique benchmark for testing both ELMS and WD atmosphere and evolutionary models. We also validate VVV 125662AB is a stable bound system despite a Galactic orbit that takes it from within the Galactic bulge to 25-35 kpc from the Galactic centre.

### v3

@@ 2 Wide binary selection
| 4757030391786232576 | GALEX J052933.2635653 | 71.2 | WD | (1) | |||
| 4757030327366948608 | 2MASS J052940266357091 | 47.92 | 70.8 | M8a | (2) | ||
| 3004346292322852096 | 2MASS J061004031031036 | 100.8 | WD | (3) | |||
| 3004346257963119872 | 2MASS J061000781031393 | 59.95 | 93.3 | M8.5b | (4) | ||
| 5863122429179888000 | VVV J125644.42620208.1 | 75.6 | DC WD | (4,5) | |||
| 5863122429178232704 | VVV J125641.09620203.8 | 18.20 | 70.4 | sdL3 | (6) | ||
| 1321738565727229056 | GALEX J155516.9+315307 | 59.8 | DA WD | (7) | |||
| 1321738561431758592 | SDSS J155517.36+315316.8 | 10.86 | 59.1 | sdM9.5a | (4) | ||
| 2056440344314369280 | IPHAS J202533.21+351509.6 | 91.1 | WD | (3) | |||
| 2056440275586819456 | 2MASS J20253351+3515035 | 7.40 | 92.2 | M8.5b | (4) |
@@
References: (1) Gaia DR3 (Gaia Collaboration et al., 2023) (2) Gentile Fusillo et al. (2021); (3) This paper; (4) DECaPS DR2 (Schlafly et al., 2018; Saydjari et al., 2023); (5) Vista VVV (Minniti et al., 2010); (6) CatWISE2020 (Marocco et al., 2021); (7) SAND atmosphere models (Alvarado et al., 2024);
@@ 4.4 The age of cool white dwarf primary
The WD companion, VVV 125662A is located at the bottom right of the WD cooling sequence in the H-R diagram (Fig. 1), consistent with being an old, low-mass WD (Raddi et al., 2022). By fitting the Gaia EDR3 photometry of this WD, Gentile Fusillo et al. (2021) estimated a K and , inferring a mass of about 0.52 M⊙. Gaia DR3 blue photometer (BP) and red photometer (RP) spectra (De Angeli et al., 2023; Montegriffo et al., 2023) are available for this source, but are rather noisy as expected for a faint object.
@@
The best-fits for our hydrogen-dominated and hydrogen-deficient models (Fig. 9) are numerically consistent, with reduced values close to unity. The hydrogen-dominated atmosphere model may be more realistic given the apparent decline in the near-infrared photometry that is not correctly captured in the models due to current uncertainties in the treatment of the H2–He collision-induced-absorption in cool WD atmospheres (Blouin et al., 2018). Our inferred atmosphere parameters of K and dex for the hydrogen-dominated model, or K and dex for the hydrogen-deficient model, yield similar masses of M⊙, and these results are compatible with those obtained by Gentile Fusillo et al. (2021).
@@ 4.5 Kinematics
We integrated the Galactic orbit of VVV 125662AB assuming a static three-component potential that is representative of the Milky Way’s baryonic and dark matter components (the MWPotential2014 of the galpy module for python; Bovy (2015) for details). We accounted for the uncertainties of the velocity components be examining both the median trajectory and 1 variants. A visual representation of the Galactic orbit and the time evolution of the cylindrical radial and vertical coordinates are shown in Fig.,11. As reported in Primeval V, this orbit is highly eccentric (), prograde ( kpc km/s), and moving radially away from the Galactic centre.
@@ 4.6 Stability of the wide binary
For VVV 125662AB, = 1.2 pc or au. Hence, this system’s projected separation of 1375 au at a distance of 75.6 pc corresponds to , well within the limit at which the Galactic tidal field is unable to disrupt the system. It is therefore logical that VVV 125662AB has survived as a bound wide binary over its 10 Gyr lifetime.
@@ 5 Conclusions
VVV 125662A is a cool, likely hydrogen-dominated WD whose total age (cooling and progenitor) of 10.5 Gyr is also consistent with halo membership. Combined, these measurements yield a consistent picture of the VVV 125662AB system as a metal-poor, wide halo binary and a unique benchmark for testing both UCD and WD atmosphere and evolutionary models. We also validate VVV 125662AB is a stable bound system despite a Galactic orbit that takes it from within the Galactic bulge to 25-35 kpc from the Galactic centre.